OpenStudy (anonymous):

(3 x 10power 6 ) (4x 10 negative power of 2 )

OpenStudy (anonymous):

I know that you combine like terms, 10^6 and 10^-2 = 10^4 and 3x4 = 12. But where does the extra power come from, i know the answer is 1.2 x 10^5. can you walk be through it?

OpenStudy (jennychan12):

just multiply. 12*10^4 12 isn't in sci. not, so move the decimal over one so it becomes 1.2*10^5

OpenStudy (anonymous):

12 is also equal to 1.2x10^1 so if you multiply 10 to 10^4 it becomes 10^5 so = 1.2*10^5